OpenStudy (anonymous):

Simplify radical expression.. square root of 10/121

OpenStudy (anonymous):

\[\sqrt{\frac{10}{121}}\]

OpenStudy (unklerhaukus):

\[\sqrt{\frac{10}{121}}=\frac{\sqrt{10}}{\sqrt{121}}\]

OpenStudy (anonymous):

yeah that's it..

OpenStudy (unklerhaukus):

now use \[11^2=121\]